An object traveling with a speed of 10 π‘š/𝑠 slides on a horizontal floor. How far will it travel before coming to rest if the coefficient of friction is 0.30?


1
Expert's answer
2021-10-10T15:54:29-0400

Given:

"\mu=0.30"

"v_0=10\:\rm m/s"

"v_f=0\:\rm m/s"


The enegry-work theorem says

"\frac{mv_f^2}{2}-\frac{mv_0^2}{2}=W=-\mu m g d"

So

"d=\frac{v_0^2}{2g\mu}=\frac{10^2}{2*9.81*0.30}=17\:\rm m"